# Does rare, noise-induced, bypass transition in plane Couette flow bypass instantons ?

### v1

@@ 2.2.3 Direct Numerical simulations of transitions
Along with the AMS computations, we perform Direct Numerical Simulations (DNS) of the build up, following a procedure we systematically use [Rolland and Simonnet, 2015, Lucente et al., 2022, Rolland, 2022]. These DNS consist in systematically repeating the first stage of AMS computations, without the additional mutation selection. Using the same definition of sets $\mathcal{A}$, $\mathcal{B}$ and hypersurface $\mathcal{C}$, we start these simulations at $\mathbf{u}=\mathbf{0}$, $p=0$ in the whole domain, let them evolve until they reach $\mathcal{C}$ and then stop them when they reach either $\mathcal{A}$ or $\mathcal{B}$. The proportion of trajectories that reach $\mathcal{B}$ before $\mathcal{A}$ provides an unbiased estimate of the crossing probability. Meanwhile the average of the duration of trajectories that reach $\mathcal{B}$ before $\mathcal{A}$ provides an unbiased estimate of the average duration of reactive trajectories. These quantities are used to validate AMS computation and detect possible problems up up to a given degree of rarity of the event. The estimates rapidly become much more costly than AMS computations as $\beta$ is increased. For very rare events other tests (mentioned in § C) are used to probe the validity of computations.
@@ 2.2.4 Methodology for bisections
We expect that in the later stages of reactive trajectories, the flow crosses the separatrix between the laminar and the turbulent basins of attractions. We will first test this by performing relaxation simulations using states taken on a reactive trajectory. In practice this consists in using states $\left(\begin{matrix}\mathbf{u}(t_{l}=l\delta t)\\ p(t_{l}=l\delta t)\end{matrix}\right)$ sampled every $\delta t>\Delta t$ along a reactive trajectory and using it as an initial condition for an unforced simulation ($\beta=0$ in equation (2)) and letting it naturally evolve, either toward the laminar flow $\mathbf{u}\rightarrow\mathbf{0}$, or toward turbulence. We can thus first detect a $t_{s}$ such that $\left(\begin{matrix}\mathbf{u}(t_{s})\\ p(t_{s})\end{matrix}\right)$ relaxes toward the laminar flow and $\left(\begin{matrix}\mathbf{u}(t_{s}+\delta t)\\ p(t_{s}+\delta t)\end{matrix}\right)$ relaxes toward the turbulent flow.
Starting from these two fields, successive shootings of a bisection procedure are then performed [Toh and Itano, 2003]. During shooting $m\geq 1$, we initialise $\eta_{0}=\frac{1}{2}$, $\eta_{\text{lam}}=1$ and $\eta_{\text{turb}}=0$. For $0\leq n\leq 36$, we run unforced simulations for $t\geq t_{m,0}$ started from
@@
    The flow laminarises, which is deemed to happen if the kinetic energy is below the threshold $E_{k}(t)\leq 0.001$. In that case we set $\eta_{n+1}=\eta_{n}-\frac{1}{2^{n+1}}$ and $\eta_{\text{lam}}=\eta_{n}$.
@@
    The flow first reaches turbulence, which is deemed to happen if the kinetic energy is above the threshold $E_{k}(t)\geq 0.07$. In that case we set $\eta_{n+1}=\eta_{n}+\frac{1}{2^{n+1}}$ and $\eta_{\text{turb}}=\eta_{n}$.
@@
$\left(\begin{matrix}\mathbf{u}_{m,-}\\ p_{m,-}\end{matrix}\right)(t=t_{m,0})=\eta_{\text{lam}}\left(\begin{matrix}\mathbf{u}_{\text{lam},m}\\ p_{\text{lam},m}\end{matrix}\right)+(1-\eta_{\text{lam}})\left(\begin{matrix}\mathbf{u}_{\text{turb},m}\\ p_{\text{turb},m}\end{matrix}\right)\,,$ (25)
@@ 3 Paths toward turbulence
In that section we describe the paths followed by the flow from the neighbourhood of the laminar state to wall turbulence under forcing, computed by means of AMS. In the first subsection 3.1, We will first describe full velocity fields on selected paths (§ 3.1.1), in the last state at the last stage (§ 3.1.2) and after bisections (§ 3.1.3). In the second subsection 3.2, we will place these observations in the statistical description of all the paths we have computed.
@@ 3.1 Velocity fields in selected examples of transition paths
We will first visualise selected paths to examine the physical mechanism of crossing from laminar to turbulent flow under forcing. The visualisations will be complemented by examination of the last state at the last stages as well as bisections started from velocity fields taken from the reactive trajectories. We will investigate whether the trajectories follow the classical instanton phenomenology, whether they cross a separatrix and in that case, whether they cross it near an unstable fixed point.
@@ 3.1.1 Description of fields on paths
We will observe the transition paths on four examples:
  • In a domain of size $L_{x}\times L_{x}=6\times 4$ at $R=500$ under divergence free forcing at high resolution (Fig. 5) using $N=20$, $N_{c}=1$, $\beta=5.5\cdot 10^{4}$, leading to $\alpha=0.35$ and $T=1.2\cdot 10^{3}$.
  • In a domain of size $L_{x}\times L_{x}=6\times 4$ at $R=500$ under a forcing on component $x$ at our standard resolution (Fig. 6), using $\beta=1.1\cdot 10^{5}$, $N=50$, $N_{c}=1$, leading to $\alpha=2.0\cdot 10^{-3}\pm 4.0\cdot 10^{-4}$ and $T=5.3\cdot 10^{5}\pm 1.8\cdot 10^{5}$.
  • In a domain of size $L_{x}\times L_{z}=18\times 12$ (Fig. 7) at $R=500$ and under a divergence free forcing at standard resolution using $N=20$, $N_{c}=1$, $\beta=6.25\cdot 10^{4}$, leading to $\alpha=0.32$ and $T=6\cdot 10^{3}$.
  • In a domain of size $L_{x}\times L_{z}=36\times 24$ (Fig. 8) at $R=500$ and under a divergence free forcing at standard resolution using $\beta=3\cdot 10^{4}$, $N=20$ and $N_{c}=4$ leading to $\alpha=0.24$ at $T=3.9\cdot 10^{3}$.
@@
Note that computations at $R=400$ in the MFU type domain and in a domain of moderate size $L_{x}\times L_{z}=12\times 8$ at $R=500$ at standard resolution were also performed, yielding similar results (not shown here). Note also that regimes of parameters leading to much smaller probability of build up $\alpha$ and much larger waiting times before build up will be presented later in this text (in this section and in § 3.2 and § 4). The same phenomenology was observed in those cases. The parameter defining the reaction coordinate for these computations are presented in appendix B.
@@
Our visualisations will be centred on the time series of kinetic energy during reactive trajectories and relaxation/bisection trajectories constructed from the reactive trajectories (Fig. 5 (f), Fig. 6 (e), Fig. 7 (g), Fig. 8 (g)). In all four cases, the time series of $E_{k}(t)$ indicate that after crossing $\mathcal{C}$, the kinetic energy first grows gradually. It does so by receiving energy in streamwise velocity tubes that show little trace of streamwise modulation and have a typical spacing slightly shorter than that of velocity streaks (Fig. 5 (c), Fig. 6 (b), Fig. 7 (b)), Fig. 8 (b)). The energy contained in the flow increases rapidly after $t\simeq 400$ in the MFU type system at higher resolution (Fig. 5 (f)), after $t\simeq 1000$ in the MFU type system forced along $\mathbf{e}_{x}$ (Fig. 6 (e)), after $t\simeq 150$ in the system of size $L_{x}\times L_{z}=18\times 12$ (Fig. 7 (g)) and after $t\simeq 200$ in the system of size $L_{x}\times L_{z}=36\times 24$ (Fig. 8 (g)). At that time, streamwise modulation of streamwise velocity, with a spanwise localisation in these larger domains, can be can be observed (Fig. 7 (d), Fig. 8 (d)). Concomitantly, streamwise vorticity with a near turbulent amplitude is also observed where $u_{x}$ is modulated (Fig. 5 (b), Fig. 7 (c), Fig. 8 (c)). The flow then has received enough energy to start the SSP on its own, even locally and finally evolve toward transitional turbulence globally. We note that this two stage process is visible in the $\log(E_{k,x}),\log(E_{k,y-z})$ plane for an example of reactive trajectory computed in a MFU type system (Fig. 11 (a)) at standard resolution for divergence free forcing using $\beta=6\cdot 10^{5}$ (leading to $\alpha=2.2\cdot 10^{-25}\pm 0.8\cdot 10^{-25}$): we first observe an increase of $E_{k,x}$, corresponding to the growth in intensity of velocity streaks, at rather low $E_{k,y-z}$ (negligible streamwise vortices). The next stage of rapid growth of kinetic energy then corresponds to the concomitant growth of $E_{k,x}$ and $E_{k,y-z}$. The same reactive trajectory leaves a less clear trace in the $I,D$ plane (Fig. 11 (b)), where both $I$ and $D$ grow in a correlated manner from $1$. We note however that $D$ fluctuates more than $I$ and that we have $D>I$ on the reactive trajectory. Being quadratic, $D$ is much more sensitive to fluctuations of velocity caused by forcing noise, which can average out in the linear $I$ (Eq. (9)). We find that $D-I\simeq 0.4\pm 0.1$ for $\beta=6\cdot 10^{6}$, corresponding to $(D-I)/R\simeq 8\cdot 10^{-4}$. Since we have $\epsilon\geq 10^{-3}$ for our control parameters, the growth of total energy is indeed driven by energy injection by the forcing, but part of that injected energy is spent in dissipation of the smaller scales generated by the noise forcing. We further checked this assertion by computing the average and standard deviation of $I-D$ on reactive trajectories in the range $3\cdot 10^{5}\leq\beta\leq 6\cdot 10^{5}$, both are affine when displayed in log-log scale as a function of $\beta$ with a $-1$ slope, indicating that they scale like $\frac{1}{\beta}$.
@@
3.1.2 Last state at the last stage
We add visualisations of the last state at the last stage to our visualisations of the reactive trajectories, following the definition of this state in section 2.2. Those states were obtained from AMS computations using the following parameters
@@ 3.1.2 Last state at the last stage
  • In a domain of size $L_{x}\times L_{x}=6\times 4$ at $R=500$ under divergence free forcing at high resolution (Fig. 9 (a,b)) using $N=400$, $N_{c}=80$, $\beta=6\cdot 10^{5}$, leading to $\alpha=2.2\cdot 10^{-25}\pm 0.8\cdot 10^{-25}$ and $T=6\cdot 10^{27}\pm 2\cdot 10^{27}$.
  • In a domain of size $L_{x}\times L_{x}=6\times 4$ at $R=500$ under a forcing on component $x$ at our standard resolution (Fig. 9 (c,d)), using $\beta=1.1\cdot 10^{5}$, $N=50$, $N_{c}=1$, leading to $\alpha=2.0\cdot 10^{-3}\pm 4.0\cdot 10^{-4}$ and $T=5.3\cdot 10^{5}\pm 1.8\cdot 10^{5}$.
  • In a domain of size $L_{x}\times L_{z}=18\times 12$ (Fig. 9 (e,f)) at $R=500$ and under a divergence free forcing at standard resolution using $N=80$, $N_{c}=16$, $\beta=8\cdot 10^{4}$, leading to $\alpha=7.7\cdot 10^{-6}$ and $T=1.8\cdot 10^{7}$.
  • In a domain of size $L_{x}\times L_{z}=36\times 24$ (Fig. 9 (g,h)) at $R=500$ and under a divergence free forcing at standard resolution using $\beta=3\cdot 10^{4}$, $N=20$ and $N_{c}=4$ leading to $\alpha=0.24$ at $T=3.9\cdot 10^{3}$.
@@
In the aforementioned figure, we include the streamwise velocity in the $y=0$ plane of this state (Fig. 9 (a,c,e,g)) and the streamwise vorticity in the $y=0$ plane of this state (Fig. 9 (b,d,f,h)), obtained form the AMS computations that yielded the reactive trajectories presented in section 3.1.1. This diagnostic indicates that at this turning point of AMS computations, streamwise elongated streamwise velocity tubes are formed and display turbulent like intensity with $\max_{\text{space}}|u_{x}|\geq 0.5$, although they present little trace of streamwise modulation in the smaller, MFU type, system (Fig. 9 (a,c)). The corresponding streamwise vorticity (Fig. 9 (b,d)) thus has a maximum intensity of $|\omega_{x}|\lesssim 0.5$, which is a fraction of what can be found when the flow is actually turbulent. Again, we note that there is no striking qualitative difference in those states obtained two distinct types of forcing. Streamwise velocity tubes are also formed in larger systems (Fig. 9 (e,g)). However, we note that in a restricted range of $z$ (taking into account periodic boundary conditions, $0\lesssim z\lesssim 3$ and $10\lesssim z\lesssim 12$ for $L_{x}\times L_{z}=18\times 12$ Fig. 9 (e), $15\lesssim z\lesssim 24$ for $L_{x}\times L_{z}=36\times 24$ Fig. 9 (g)), wall turbulence like streamwise modulation of $u_{x}$ is present. We note that streamwise vorticity of turbulent like intensity $\omega_{x}=\mathcal{O}(1)$ and with turbulent like scale is now only found in a similar range of $z$ (taking into account periodic boundary conditions, $10\lesssim z\lesssim 12$ and to a lesser extent $0\lesssim z\lesssim 3$ for $L_{x}\times L_{z}=18\times 12$ Fig. 9 (f), $18\lesssim z\lesssim 24$ for $L_{x}\times L_{z}=36\times 24$ Fig. 9 (h)). Vorticity is more intense in the larger system. This further indicates that while the reactive trajectory corresponds to a response to the forcing through the development of streamwise velocity tube, the flow only reach a turning point when locally, the tubes become sufficiently intense and display intense enough shear between one another to trigger the formation of streamwise modulation and corresponding streamwise vortices. Once this has happened, this streamwise modulation and said vortices will contaminate nearby regions in $z$ until they extend to the whole flow.
@@
Figure 9: From the AMS computation in the system of size $L_{x}\times L_{z}=6\times 4$ at $R=500$ and $\beta=5.5\cdot 10^{5}$ at high resolution, (a) colour levels of the streamwise velocity in the plane $y=0$ on the last state at the last stage for the AMS computation, (b) colour levels of the streamwise vorticity in the plane $y=0$ on the last state at the last stage for this AMS computation. From the AMS computation in the system of size $L_{x}\times L_{z}=6\times 4$ at $R=500$ and $\beta=6\cdot 10^{5}$ at sandard resolution, (c) colour levels of the streamwise velocity in the plane $y=0$ on the last state at the last stage for the AMS computation, (d) colour levels of the streamwise vorticity in the plane $y=0$ on the last state at the last stage for this AMS computation. From the AMS computation in the system of size $L_{x}\times L_{z}=18\times 12$ at $R=500$, and $\beta=8\cdot 10^{4}$, (e) colour levels of the streamwise velocity in the plane $y=0$ on the last state at the last stage for the AMS computation, (f) colour levels of the streamwise vorticity in the plane $y=0$ on the last state at the last stage for this AMS computation. From the AMS computation in the system of size $L_{x}\times L_{z}=36\times 24$ at $R=500$, at $\beta=3\cdot 10^{4}$, (g) Colour levels of the streamwise velocity in the plane $y=0$ on the last state at the last stage for this AMS computation, (h) Colour levels of the streamwise vorticity in the plane $y=0$ on the last state at the last stage for this AMS computation.

### v2

@@ 2.2.3 Direct Numerical simulations of transitions
Along with the AMS computations, we perform Direct Numerical Simulations (DNS) of the build up, following a procedure we systematically use Rolland and Simonnet, (2015); Lucente et al., (2022); Rolland, (2022). These DNS consist in systematically repeating the first stage of AMS computations, without the additional mutation selection. Using the same definition of sets $\mathcal{A}$, $\mathcal{B}$ and hypersurface $\mathcal{C}$, we start these simulations at $\mathbf{u}=\mathbf{0}$, $p=0$ in the whole domain, let them evolve until they reach $\mathcal{C}$ and then stop them when they reach either $\mathcal{A}$ or $\mathcal{B}$. The proportion of trajectories that reach $\mathcal{B}$ before $\mathcal{A}$ provides an unbiased estimate of the crossing probability. Meanwhile the average of the duration of trajectories that reach $\mathcal{B}$ before $\mathcal{A}$ provides an unbiased estimate of the average duration of reactive trajectories. These quantities are used to validate AMS computation and detect possible problems up up to a given degree of rarity of the event. The estimates rapidly become much more costly than AMS computations as $\beta$ is increased, because the proportion of trajectories that go straight back to $\mathcal{A}$ becomes almost $1$, in agreement with the description of the forced flow as having two narrow mulistable states. For very rare events other tests (mentioned in § D) are used to probe the validity of computations.
@@ 2.2.4 Methodology for bisections
We expect that in the relaxation path of noise induced reactive trajectories, the flow crosses the separatrix between the laminar and the turbulent basins of attractions of the deterministic system. This supposition would be in line with the statistical physics description of bistability under a vanishing noise. We will first test this by performing relaxation simulations using states taken on a reactive trajectory. In practice this consists in using states $\left(\begin{matrix}\mathbf{u}(t_{l}=l\delta t)\\ p(t_{l}=l\delta t)\end{matrix}\right)$ sampled every $\delta t>\Delta t$ along a reactive trajectory and using it as an initial condition for an unforced simulation ($\beta=0$ in equation (2)) and letting it naturally evolve, either toward the laminar flow $\mathbf{u}\rightarrow\mathbf{0}$, or toward turbulence. We can thus first detect a $t_{s}$ such that $\left(\begin{matrix}\mathbf{u}(t_{s})\\ p(t_{s})\end{matrix}\right)$ relaxes toward the laminar flow and $\left(\begin{matrix}\mathbf{u}(t_{s}+\delta t)\\ p(t_{s}+\delta t)\end{matrix}\right)$ relaxes toward the turbulent flow.
Starting from these two fields, successive shootings of a bisection procedure are then performed Toh and Itano, (2003). During shooting $m\geq 1$, we initialise $\eta_{0}=\frac{1}{2}$, $\eta_{\text{lam}}=1$ and $\eta_{\text{turb}}=0$. For $0\leq n\leq 36$, we run unforced simulations for $t\geq t_{m,0}$ started from
@@
    The flow laminarises, which is deemed to happen if the kinetic energy is below the threshold $E_{k}(t)\leq 0.001$. This threshold is chosen so that the flow has underwent extensive viscous decay when $E_{k}(t)=0.001$ is reached, so that the relaxation toward the saddle is not interrupted and no saddle is excluded. This plays a more important role in the largest domain of size $L_{x}\times L_{z}=36\times 24$ where smaller values of $E_{k}(t)$ are found because non trivial flow occupies a smaller area in the domain. In that case we set $\eta_{n+1}=\eta_{n}-\frac{1}{2^{n+1}}$ and $\eta_{\text{lam}}=\eta_{n}$.
@@
    The flow first reaches turbulence, which is deemed to happen if the kinetic energy is above the threshold $E_{k}(t)\geq 0.07$. This threshold is chosen large enough so that no saddle is mistaken for turbulent flow, for all domain sizes, but not so large that it would not be reached by wall turbulence. This plays a more important role in the smallest domain of size $L_{x}\times L_{z}=36\times 24$ where the saddle has larger values of $E_{k}(t)$ than for other domain sizes. In that case we set $\eta_{n+1}=\eta_{n}+\frac{1}{2^{n+1}}$ and $\eta_{\text{turb}}=\eta_{n}$.
@@
$\left(\begin{matrix}\mathbf{u}_{m,-}\\ p_{m,-}\end{matrix}\right)(t=t_{m,0})=\eta_{\text{lam}}\left(\begin{matrix}\mathbf{u}_{\text{lam},m}\\ p_{\text{lam},m}\end{matrix}\right)+(1-\eta_{\text{lam}})\left(\begin{matrix}\mathbf{u}_{\text{turb},m}\\ p_{\text{turb},m}\end{matrix}\right)\,,$
@@ 3 Paths toward turbulence
In that section we describe the paths followed by the flow from the neighbourhood of the laminar state to wall turbulence under forcing, computed by means of AMS. In the first subsection 3.1, We will first describe full velocity fields on selected paths (§ 3.1.1), in the last state at the last step (§ 3.1.2) and after bisections (§ 3.1.3). In the second subsection 3.2, we will place these observations in the statistical description of all the paths we have computed.
@@ 3.1 Velocity fields in selected examples of transition paths
We will first visualise selected paths to examine the physical mechanism of crossing from laminar to turbulent flow under forcing. The visualisations will be complemented by examination of the last states at the last step as well as bisections started from velocity fields taken from the reactive trajectories. We will investigate whether the trajectories follow the classical instanton phenomenology, whether they cross a separatrix and in that case, whether they cross it near an unstable fixed point.
@@ 3.1.1 Description of fields on paths
We will observe the transition paths on four examples encompassing the three domain sizes considered in this article (Fig. 567 and 8). We list the property of the AMS computations, the parameter $\beta$ controlling the inverse of the variance of the forcing (and the energy injection rate), the estimated probability of transition and mean first passage time in table 2. Note that computations at $R=400$ in the MFU type domain and in a domain of moderate size $L_{x}\times L_{z}=12\times 8$ at $R=500$ at standard resolution were also performed, yielding similar results (not shown here). Note also that regimes of parameters leading to much smaller probability of build up $\alpha$ and much larger waiting times before build up will be presented later in this text (in this section and in § 3.2 and § 4). The same phenomenology was observed in those cases. The parameter defining the reaction coordinate for these computations are presented in appendix C.
$L_{x}\times L_{z}$ resolution $N$ $N_{c}$ $\beta$ $\alpha$ $T$ forcing type Figure $6\times 4$ high 20 1 $5.5\cdot 10^{4}$ divergence free 0.35 $1.2\cdot 10^{3}$ Fig. 5 $6\times 4$ standard 50 1 $1.1\cdot 10^{5}$ on $u_{x}$ $2.0\cdot 10^{-3}\pm 4.0\cdot 10^{-4}$ $5.3\cdot 10^{5}\pm 1.8\cdot 10^{5}$ Fig. 6 $18\times 12$ 20 1 $6.25\cdot 10^{4}$ divergence free $\alpha=0.32$ $T=6\cdot 10^{3}$ Fig. 7 $36\times 24$ 20 4 $3\cdot 10^{4}$ divergence free $\alpha=0.24$ $T=3.9\cdot 10^{3}$ Fig. 8
Table 2: List of the parameters used in AMS computations of transition paths in domains of several sizes $L_{x}\times L_{z}$ at Reynolds number $R=500$ presented in figures 567 and 8, and resulting properties. We indicate the resolution type (when relevant), the parameter $\beta$ controlling the inverse of the variance of the forcing (details on forcing spectra are given in appendix A), along with the type of forcing, the number of clones used in AMS computation $N$, as well as the number of clones removed at each step of the algorithm $N_{c}$. The estimated probability of build up $\alpha$ and mean first passage time before build up $T$ are given, along with the figure presenting the path.
@@
Our visualisations will be centred on the time series of kinetic energy during reactive trajectories and relaxation/bisection trajectories constructed from the reactive trajectories (Fig. 5 (f), Fig. 6 (e), Fig. 7 (g), Fig. 8 (g)). In all four cases, the time series of $E_{k}(t)$ indicate that after crossing $\mathcal{C}$, the kinetic energy first grows gradually. It does so by receiving energy in streamwise velocity tubes that show little trace of streamwise modulation and have a typical spacing slightly shorter than that of velocity streaks (Fig. 5 (c), Fig. 6 (b), Fig. 7 (b)), Fig. 8 (b)). The energy contained in the flow increases rapidly after $t\simeq 400$ in the MFU type system at higher resolution (Fig. 5 (f)), after $t\simeq 1000$ in the MFU type system forced along $\mathbf{e}_{x}$ (Fig. 6 (e)), after $t\simeq 150$ in the system of size $L_{x}\times L_{z}=18\times 12$ (Fig. 7 (g)) and after $t\simeq 200$ in the system of size $L_{x}\times L_{z}=36\times 24$ (Fig. 8 (g)). At that time, streamwise modulation of streamwise velocity, with a spanwise localisation in these larger domains, can be can be observed (Fig. 7 (d), Fig. 8 (d)). Concomitantly, streamwise vorticity with a near turbulent amplitude is also observed where $u_{x}$ is modulated (Fig. 5 (b), Fig. 7 (c), Fig. 8 (c)). The flow then has received enough energy to start the SSP on its own, even locally and finally evolve toward transitional turbulence globally. We note that this two stage process is visible in the $\log(E_{k,x}),\log(E_{k,y-z})$ plane for an example of reactive trajectory computed in a MFU type system (Fig. 11 (a)) at standard resolution for divergence free forcing using $\beta=6\cdot 10^{5}$ (leading to $\alpha=2.2\cdot 10^{-25}\pm 0.8\cdot 10^{-25}$): we first observe an increase of $E_{k,x}$, corresponding to the growth in intensity of velocity streaks, at rather low $E_{k,y-z}$ (negligible streamwise vortices). The next stage of rapid growth of kinetic energy then corresponds to the concomitant growth of $E_{k,x}$ and $E_{k,y-z}$. The same reactive trajectory leaves a less clear trace in the plane formed by the energy injection $I$ and the dissipation $D$ (Fig. 11 (b)), where both $I$ and $D$ grow in a correlated manner from $1$. We note however that the dissipation $D$ fluctuates more than the energy injection $I$ and that we have $D>I$ on the reactive trajectory. Being quadratic, $D$ is much more sensitive to fluctuations of velocity caused by forcing noise, which can average out in the linear $I$ (Eq. (10)). We find that $D-I\simeq 0.4\pm 0.1$ for $\beta=6\cdot 10^{6}$, corresponding to $(D-I)/R\simeq 8\cdot 10^{-4}$. Since we have $\epsilon\geq 10^{-3}$ for our control parameters, the growth of total energy is indeed driven by energy injection by the forcing, but part of that injected energy is spent in dissipation of the smaller scales generated by the noise forcing. We further checked this assertion by computing the average and standard deviation of $I-D$ on reactive trajectories in the range $3\cdot 10^{5}\leq\beta\leq 6\cdot 10^{5}$, both are affine when displayed in log-log scale as a function of $\beta$ with a $-1$ slope, indicating that they scale like $\frac{1}{\beta}$.
@@
3.1.2 Last state at the last step
@@
$L_{x}\times L_{z}$ resolution $N$ $N_{c}$ $\beta$ forcing type $\alpha$ $T$ figure $6\times 4$ high 400 80 $6\cdot 10^{5}$ divergence free $2.2\cdot 10^{-25}\pm 0.8\cdot 10^{-25}$ $6\cdot 10^{27}\pm 2\cdot 10^{27}$ Fig. 9 (a,b) $6\times 4$ standard 50 1 $1.1\cdot 10^{5}$ on $u_{x}$ $2.0\cdot 10^{-3}\pm 4.0\cdot 10^{-4}$ $5.3\cdot 10^{5}\pm 1.8\cdot 10^{5}$ Fig. 9 (c,d) $18\times 12$ 80 16 $8\cdot 10^{4}$ divergence free $7.7\cdot 10^{-6}$ $1.8\cdot 10^{7}$ Fig. 9 (e,f) $36\times 24$ 20 4 $3\cdot 10^{4}$ divergence free $0.24$ $3.9\cdot 10^{3}$ Fig. 9 (g,h)
Table 3: Table indicating the parameters used in AMS computations where the last states at the last step presented in figure 9 were computed. This indicates the size of the domains $L_{x}\times L_{z}$ where the computations were performed at Reynolds number $R=500$ and the resolution type (when relevant). We give the number of clones $N$ used in AMS computations along with the number of clones removed at each step $N_{c}$, the parameter $\beta$ controlling the inverse of the variance of the forcing and the energy injection rate, as well as the forcing type. The estimated properties (probability of transition $\alpha$ and mean first passage time $T$) of the transition paths from which the states are extracted are provided, along with the label of the figure where the states are shown.
@@
We add visualisations of the last state at the last step to our visualisations of the reactive trajectories, following the definition of this state in section 2.2. Those states were obtained from AMS computations using parameters given in table 3 for all three of our domain sizes, both at standard and high resolution and for both forcing type in MFU type domains. In figure 9, we include the streamwise velocity in the $y=0$ plane of this state (Fig. 9 (a,c,e,g)) and the streamwise vorticity in the $y=0$ plane of this state (Fig. 9 (b,d,f,h)), obtained form the AMS computations that yielded the reactive trajectories presented in section 3.1.1. This diagnostic indicates that at this turning point of AMS computations, streamwise elongated streamwise velocity tubes are formed and display turbulent like intensity with $\max_{\text{space}}|u_{x}|\geq 0.5$, although they present little trace of streamwise modulation in the smaller, MFU type, system (Fig. 9 (a,c)). The corresponding streamwise vorticity (Fig. 9 (b,d)) thus has a maximum intensity of $|\omega_{x}|\lesssim 0.5$, which is a fraction of what can be found when the flow is actually turbulent. Again, we note that there is no striking qualitative difference in those states obtained two distinct types of forcing. Streamwise velocity tubes are also formed in larger systems (Fig. 9 (e,g)). However, we note that in a restricted range of $z$ (taking into account periodic boundary conditions, $0\lesssim z\lesssim 3$ and $10\lesssim z\lesssim 12$ for $L_{x}\times L_{z}=18\times 12$ Fig. 9 (e), $15\lesssim z\lesssim 24$ for $L_{x}\times L_{z}=36\times 24$ Fig. 9 (g)), wall turbulence like streamwise modulation of $u_{x}$ is present. We note that streamwise vorticity of turbulent like intensity $\omega_{x}=\mathcal{O}(1)$ and with turbulent like scale is now only found in a similar range of $z$ (taking into account periodic boundary conditions, $10\lesssim z\lesssim 12$ and to a lesser extent $0\lesssim z\lesssim 3$ for $L_{x}\times L_{z}=18\times 12$ Fig. 9 (f), $18\lesssim z\lesssim 24$ for $L_{x}\times L_{z}=36\times 24$ Fig. 9 (h)). Vorticity is more intense in the larger system. This further indicates that while the reactive trajectory corresponds to a response to the forcing through the development of streamwise velocity tube, the flow only reach a turning point when locally, the tubes become sufficiently intense and display intense enough shear between one another to trigger the formation of streamwise modulation and corresponding streamwise vortices. Once this has happened, this streamwise modulation and said vortices will contaminate nearby regions in $z$ until they extend to the whole flow.
@@
Figure 9: From the AMS computation in the system of size $L_{x}\times L_{z}=6\times 4$ at $R=500$ and $\beta=5.5\cdot 10^{5}$ at high resolution, (a) colour levels of the streamwise velocity in the plane $y=0$ on the last state at the last step for the AMS computation, (b) colour levels of the streamwise vorticity in the plane $y=0$ on the last state at the last step for this AMS computation. From the AMS computation in the system of size $L_{x}\times L_{z}=6\times 4$ at $R=500$ and $\beta=6\cdot 10^{5}$ at sandard resolution, (c) colour levels of the streamwise velocity in the plane $y=0$ on the last state at the last step for the AMS computation, (d) colour levels of the streamwise vorticity in the plane $y=0$ on the last state at the last step for this AMS computation. From the AMS computation in the system of size $L_{x}\times L_{z}=18\times 12$ at $R=500$, and $\beta=8\cdot 10^{4}$, (e) colour levels of the streamwise velocity in the plane $y=0$ on the last state at the last step for the AMS computation, (f) colour levels of the streamwise vorticity in the plane $y=0$ on the last state at the last step for this AMS computation. From the AMS computation in the system of size $L_{x}\times L_{z}=36\times 24$ at $R=500$, at $\beta=3\cdot 10^{4}$, (g) Colour levels of the streamwise velocity in the plane $y=0$ on the last state at the last step for this AMS computation, (h) Colour levels of the streamwise vorticity in the plane $y=0$ on the last state at the last step for this AMS computation.